### v1

# Debt, Growth, and the Carbon Lock-In

Abstract
Despite decades of climate policy and technological progress, global CO2 emissions continue to rise while natural carbon sinks weaken. Here we provide a structural explanation for this persistent overshoot by identifying a link between financial leverage and emissions. We develop a macro-financial model that connects credit dynamics, economic growth, and cumulative carbon emissions. The model extends Kelly’s information-theoretic growth framework to incorporate debt, bankruptcy risk, and emissions trajectories, showing that borrowing amplifies GDP growth and cumulative emissions, thereby locking economies into high-carbon pathways even as energy efficiency improves.
This approach complements traditional Integrated Assessment Models (IAMs) by explicitly linking financial dynamics to cumulative emissions, enabling assessment of financial policy alignment with net-zero pathways. We show that debt increases short-term production gains but at the cost of higher bankruptcy risk and higher cumulative emissions. The model reveals a double constraint: financially, debt requires growth to be repaid; ecologically, growth requires energy, which in turn entails emissions. When intrinsic growth falls below the interest rate, solvency collapses, identifying an optimal leverage frontier beyond which additional debt no longer fuels real wealth and only increases climatic and financial risk.
We calibrate the model using multi-decade macroeconomic and emissions data for the United States, China, France, and Denmark. Across all cases, debt levels, cumulative GDP, and cumulative emissions remain tightly correlated, challenging the feasibility of growth–emissions decoupling and indicating that net-zero targets cannot be achieved without aligning credit allocation with decarbonisation objectives.
In our current financial system, credit expansion amplifies GDP growth and associated emissions, thereby locking economies into higher cumulative carbon trajectories, even as energy efficiency improves through innovation.
keywords:
Leverage-driven growth, Default risk, Debt-emissions feedback, Carbon lock-in, Double materiality, Ecological macroeconomic modeling, Climate finance
Introduction
Debt is a social construct [1, 2], which may appear, for this reason, disconnected from the physical reality. However, debt has real-world implications because borrowed money facilitates the growth of businesses by enabling the consumption of energy and resources, which may be available in finite amounts. Alternatively, it can also cause financial instabilities due to leverage cycles  [3]. Thus, debt brings consumption and production forward by enabling energy investment in resource-intensive sectors, including marginal resource extraction and high-cost infrastructure [2].
In fact, debt creates new feedback loops such as the one represented in Fig. 1, in which credit expansion stimulates growth, which in turn requires additional energy (often directed towards carbon-intensive sectors) and results in carbon emissions, despite ongoing efforts towards decarbonization. In the literature, such a feedback loop is related to the so-called ”macro-economic rebound effect” when investment in clean energy drives productivity increases, which drives growth, thereby also increasing emissions [4].
Figure 1: The debt-carbon spiral: leverage enhances economic growth, but it also increases emissions and default risk, thereby establishing a feedback loop that locks in carbon-intensive development and jeopardizes climate targets and financial stability. In this scheme, the letters (R) and (B) represent reinforcing and balancing loops, respectively. The primary reinforcing loop (R1) illustrates how the expansion of credit and leverage stimulates GDP growth, leading to increased energy consumption and CO2 emissions. A secondary reinforcing loop (R2) describe the necessity of additional borrowing to maintain the same level of resource extraction as the energy return on investment (EROI) declines and the costs relative to climate change damages rise [5, 6]. This, in turn, leads to more aggressive fiscal and credit policies that further intensify debt-financed growth. Two balancing loops counteract these self-propelling cycles: (B1) captures the negative economic effects of environmental externalities, such as pollution, climate-related losses, and reduced productivity, that constrain growth [6], while (B2) captures technological and structural decarbonization mechanisms that reduce the carbon intensity of energy use and improve efficiency. Feedback that are qualitatively supported in the literature but are not explicitly quantified in our model are indicated by dashed lines.
Similar feedback loops exist in other fields such as biology. In biology, living cells adjust to environmental constraints by employing two complementary strategies: (i) decreasing their metabolic rate to consume less energy per unit time, or (ii) reconfiguring their metabolic pathways to optimize the use of available energy or alternative sources [7, 8, 9, 10]. In a similar manner, economies that are confronted with resource constraints have the option to either (i) decrease the energy required per unit of output, thereby reducing the carbon intensity of GDP, or (ii) transition their energy mix to low-carbon sources [11]. An important limitation to this analogy between biology and economy, however, is that there is no biological equivalent of debt. Debt is thus inherent to financial systems; it facilitates growth on the promise of future expected gains in a manner that biological systems cannot, thereby establishing new mechanisms of instability due to an increase of emissions, resource depletion, socio-economic destabilization and and ecosystem degradation [12, 13].
Motivated by this observation, we study in this paper the link between finance and climate change, focusing particularly on the role of debt. Another motivation for this study is that standard Integrated Assessment Models (IAMs), which typically represent emissions as a function of exogenous GDP growth  [14, 15], do not model the financial system nor investors’ decisions. Thus, the feedback loop represented in Fig. 1 that couples the financial system and mitigation pathways is typically not taken into account by IAMs. Several authors have identified this shortcoming and have criticized IAMs for their lack of financial system [16, 17].
To model the financial system, we rely on Kelly’s gambling model [18], because this model offers a general theory of money, which is not based on notions of economic equilibrium. In 1956, Kelly found a deep connection between information theory and gambling in horse races. This work then became highly influential in investment science [19, 20] and in evolutionary biology [21]. Kelly’s optimal strategy defines an optimal capital allocation by maximizing the expected logarithmic growth rate. The model demonstrates that diversification of assets enables a mitigation of risk from environmental fluctuations, a goal that is also central in biology and ecology and is related to concepts of bet-hedging [11, 21, 22]. In this analogy, the growth of the gambler’s capital is equivalent to the growth of the population, and the risk of bankruptcy is comparable to the extinction of the population. Despite diversification, Kelly’s optimal strategy is known to be risky and this observation hints at a general trade-off between the risk of bankruptcy and the average long-term growth rate of the capital [23, 24].
The original Kelly’s model, however, does not include the possibility of bankruptcy nor the possibility of borrowing(debt). To model these effects, we extend Kelly’s framework to include default dynamics under compounding interest obligations. Once the borrower has taken out a loan, we assume that the debt accumulates with an interest rate at each time step. When the better pays back the debt, the capital is reset at a lower value. The evolution of the capital can thus be compared to a biased random walk with resetting [25, 26, 27], or to a coupled dynamics of money and anti-money variables [28].
In this study, we will use this model as a basis to describe the critical role that credit expansion plays in driving economic growth and cumulative CO2 emissions. Besides quantifying this effect, our stochastic framework also establishes an optimal leverage frontier: the probability of long-term solvency approaches zero when the intrinsic growth rate (growth that originates within an economy, fuelled by factors such as productivity improvements) falls below the effective interest rate, suggesting that structural financial instability is the ultimate effect of excessive borrowing. In this work, we show that credit is the primary driver of several major economies and that this comes at the cost of increasing global dissipation in the form of emissions  [29, 17]. We used country-level economic data to predict the cumulative emissions per capita for these countries. Cumulative CO2 emissions, rather than CO2 emissions, are a key quantity that captures the effect of economic activities in increasing global warming. We demonstrate that to fuel economic growth, credit is required, which has a cost in terms of CO2 emissions.
Our contribution complements standard Integrated Assessment Models (IAMs), which typically do not aim to reproduce past emissions, GDP co-evolution, or assess the causal role of credit in driving the rise in emissions. In contrast, our model uses empirical fiscal, macroeconomic, and climate data to explicitly link debt-financed growth with cumulative emissions, thereby offering a plausible explanation for the persistent proportionality between cumulative GDP and emissions, thus assessing the responsibility for cumulative carbon emissions even in countries with fast decarbonization trajectories. Our approach aligns more closely with recent works in macro-finance that analyze climate-related financial risk [17], and with studies in ecological macroeconomics [30, 31] and system dynamics [32] that attempt to take into account socio-environmental feedback loops similar to the debt-carbon spiral presented in this paper (Fig. 1).
Model
We present now our extension of Kelly’s model, and we explain in SI how known models of public debt  [33] can be mapped on this framework in sections 1.D and 1.E. Let us assume that at each time step $\tau$ (we use discrete time), the gambler can borrow an amount $B_{\tau}$ with interest rate $\rho_{\tau}$ (real interest rate accounting for inflation [34]). The interest rate $\rho_{\tau}$ is the factor by which the amount borrowed is multiplied at each time step, in other words, the debt due to the amount borrowed at time $\tau^{\prime}$, is multiplied by $\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}$ if it has not been paid back at time $\tau$. In general, the interest rate is larger than 11 ($\rho>1$), so that the debt keeps increasing, and the borrower ends up paying more than the amount borrowed. Then the gross capital $C_{\tau+1}$, and debt $D_{\tau+1}$ after one step are (until bankruptcy occurs if $C_{\tau+1}\leq 0$):
$C_{0}$ $=B_{0}+C_{\text{ini}}=LC_{\text{ini}},$
$C_{\tau+1}$ $=\gamma_{\tau}C_{\tau}+B_{\tau+1}-\left(\sum_{\tau^{\prime}=0}^{\tau}\delta_{\tau,t_{p}(\tau^{\prime})}\right)D_{\tau+1},$
$D_{0}$ $=B_{0}=(L-1)C_{\text{ini}},$
$D_{\tau}$ $=\sum_{\tau^{\prime}=0}^{\tau-1}\Theta(t_{p}(\tau^{\prime})-\tau)\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}B_{\tau^{\prime}},$
where $\gamma_{\tau}$ (also accounting for inflation [34]) is the stochastic instantaneous growth rate and depends on time (in a classical Kelly’s model for horse races this factor would equal the product of the bets and odds on the winning horse [23, 24]). The first term in the first equation represents the standard evolution in Kelly’s model, namely the contribution of gambling at time $\tau$: if $\gamma_{\tau}>1$, the gross capital increases provided that there is no borrowing or payback at this time. The second term represents the amount borrowed at time $\tau$; it can be 0 if the borrower does not borrow at this step, or a positive value. The third term represents the possible payback of the debt (debt clearing) at time $\tau+1$, and will either be equal to 0 (no payback), or to $D_{\tau+1}$. $C_{\text{ini}}$ is the amount invested initially by the gambler, and $B_{0}$ is the initial amount borrowed. As usual, the leverage is the ratio of the asset value to the cash needed to purchase it, namely $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$ [3].
In addition, $t_{p}(\tau^{\prime})$ indicates the time when the amount borrowed at $\tau^{\prime}$ is paid back. The Heaviside function $\Theta$ function allows us to know if the payback has occurred at time $\tau$ (indeed $\Theta(t_{p}(\tau^{\prime})-\tau)=0$ if $\tau>t_{p}$, meaning that the debt is erased after it has been payed back). If a payback happens at time $\tau$, it means that $\exists\tau^{\prime},\tau=t_{p}(\tau^{\prime})$, and thus $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=1$, otherwise $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=0$.
Without debt, the gross capital cannot hit 0 in Kelly’s model, and bankruptcy never occurs. In our updated version of this model, the capital can become negative if the debt exceeds the capital at the time the debt is repaid. Therefore solvency means that the debt is always paid back at a time where the gross capital is larger ($\gamma_{\tau}C_{\tau}>D_{\tau+1}$). We call $\gamma_{\tau}C_{\tau}-D_{\tau+1}$ the net capital, whose sign indicates if bankruptcy will happen in case of payback. Empirically, the estimated leverage $L_{t}$ remains fairly stable within each economy, varying by only a few percent across years (as shown in Fig. S1 in the SI). This supports the hypothesis of quasi-stationary financial amplification of growth. In the following sections, we outline two different strategies that represent two distinct real-world debt-financed growth regimes.
Finite leverage — initial investment (strategy A)
We assume that the gambler initially borrows money ($B_{0}>0$) but never borrows afterwards, $\forall\tau\geq 1,B_{\tau}=0$. We refer to this strategy as strategy A, which is akin to an individual taking out a loan to buy a house. In practice, the gambler can borrow $B_{0}$ and invest $C_{\text{ini}}$ from its personal capital, and the leverage is then $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model with an initial investment of $B_{0}+C_{\text{ini}}$, and the debt evolves with an interest rate $\rho$. At the time of payback, the capital can become negative, which could not happen within the purely multiplicative model. Bankruptcy occurs if $C_{t_{p}}$ is negative, or equivalently if $\prod_{\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}<1-\frac{1}{L}$, meaning that the larger $L$ is, the more likely the bankruptcy. Under the assumption of no bankruptcy, we can also define the long-term growth rate $W=\mathbb{E}\left({\ln\left(\gamma\right)}\right)$. Note that due to the concavity of the logarithm, we have generally for any probability distribution of $\gamma$,
$\ln(E)\geq W,$ (1)
with $E=\mathbb{E}\left({\gamma}\right)$. However, there is a non-zero probability of hitting bankruptcy when the debt is paid back (at $t_{p}$).
Persistent leverage — borrowing at each time step (strategy B)
Let us instead assume that the gambler borrows money at each time step, with leverage $L_{\tau}$ (now time-dependent), which we refer to as strategy B. This strategy is rather that of a state taking yearly on credit to ensure its budget balance. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model, with a modified stochastic growth rate, and the debt still evolves with an interest rate $\rho$. The dynamics are then (except for the time of payback)
$\begin{split}C_{\tau+1}&=\left(\gamma_{\tau}+(L_{\tau}-1)\right)C_{\tau},\\ D_{\tau+1}&=\rho D_{\tau}+(L_{\tau}-1)C_{\tau}.\end{split}$ (2)
Now, in addition to the long-term growth rate $W$ defined as before, one needs to also consider the short-term growth rate $\omega(t_{p})$ (defined in SI 1. B). We can separate the value added from production, and that which is due to borrowing in the evolution equations, and define the stochastic intrinsic growth rate as the growth rate to which the part due to borrowing [33] is subtracted:
$r(\tau)=\frac{C_{\tau+1}-C_{\tau}}{C_{\tau}}-(L_{\tau}-1)=\gamma_{\tau}-1.$ (3)
Results
Lock-in of carbon emissions
Now, we scale up the above financial model with strategy B to describe the wealth and the carbon emissions of a country. Let $C_{\tau}$ represents the economic wealth in the year $\tau$ of that country, and let us assume that an external bank lends money each year to that country. As a result, its debt, encapsulated by the variable $D_{\tau}$ increases. Let us also assume that the leverage is constant within the time frame considered. This assumption is justified by the analysis of economic data shown in Fig. S1.
In practice, the GDP during a given period is often used as a measure of economic wealth and has also been criticized for this reason [35, 36]. It consists of summing the monetary values of all products created; therefore, the added growth product between time $\tau$ and $\tau+1$ is $C_{\tau+1}-C_{\tau}$. This capital will be invested in activities that require some energy and produce carbon emissions at time $\tau$, described by the function $\epsilon(\tau,C_{\tau})$. This should be an increasing function of capital, and thus its average should increase with leverage (i.e., the ratio of total money invested to money borrowed). We ultimately have that $\epsilon$ is an increasing function of the capital $C_{\tau}$, which is observed in practice [37].
A key quantity is the carbon intensity of the economy $I(\tau)$ [38], representing the carbon cost of one unit of $GDP$. We connect the emissions to the carbon intensity of the economy thanks to the Kaya relationship [38, 39, 14, 40]:
$\epsilon(\tau,C_{\tau})=C_{\tau}\times I(\tau).$ (4)
The parameter $I(\tau)$ is typically a decreasing function of $\tau$ [38, 14, 40] because energy mixes tend to incorporate more renewable energies over time. Note that $I(\tau)$ can decrease for two main reasons, as illustrated in Fig.1, either the energy cost of $1USD$ decreases, meaning that money is used in less energy-intensive activities, or the carbon cost of energy decreases, meaning that the energy mix of a country includes ”greener” energies. However, this does not mean that cumulative emissions are reduced, as we will see.
Let us now examine the lock-in effect arising from path-dependent cumulative emissions. Public debt can be a powerful tool for investing in the transition, but it can also be used to pursue carbon-intensive economic activities; the difference between the two is encapsulated in the concept of carbon intensity. As this quantity changes with time [38], the cumulative emissions corresponding to a particular trajectory for a country’s GDP will also depend on the trajectory of the carbon intensity of this economy. We can measure the lock-in effect using the path-dependent intensity:
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)=\frac{\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})C_{\tau^{\prime}}}{\sum_{\tau^{\prime}=0}^{\tau}C_{\tau^{\prime}}},$ (5)
which compares cumulative carbon emissions to cumulative GDP over a given period. This quantity represents the cumulative carbon emissions for a given amount of money invested. Contrary to carbon intensities, this variable is path-dependent, similarly to heat and work in thermodynamics [41]. This quantity allows for assessing the actual responsibility of countries for cumulative carbon emissions over a given period, irrespective of recent decarbonization efforts captured by carbon intensity.
We show in SI 2.C that for any strictly decreasing carbon intensities $I(\tau)$, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ is also a decreasing function of time. We then study the influence of the decarbonization rate (or efficiency of innovations) in SI 2.C and conclude that if $I(\tau)$ is decreasing linearly, the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ stabilizes to a finite value. In general, the difference goes like $(dI/d\tau)/(e^{r}-1)$ at large times. If it decreases faster than linearly, the difference tends to $\infty$, whereas if it decreases more slowly than linearly, the difference first increases and then tends to 0. In practice, it means that the faster decarbonization occurs, the more early carbon-intensive years will matter in the path-dependent intensity. On the contrary, for slow decarbonization, recent years matter more in the path-dependent intensity, making it closer to the carbon intensity.
The larger the path-dependent intensity, the greater the lock-in effect, because we cannot reset $CO_{2}$ emissions to 0 every year. We show path-dependent intensities, as well as intensities in Fig.2. We observe that the orders of magnitude differ widely among countries. We also observe that even if carbon intensities decrease over time, path-dependent intensities remain practically unchanged over a 20-year period. This means that cumulative emissions are largely proportional to cumulative GDP, despite efforts to reduce the annual carbon intensity.
The difference between annual carbon intensity and path-dependent intensity is detailed in the SI. In Fig.2C, we show the correlation between cumulative emissions per capita (since 1998) and cumulative GDP (since 1998) for each country. What we find is that for a carbon intensity decreasing with time and a growing economy, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ is larger than $I(\tau)$. In particular, the difference between the two depends on the long-term growth rate $W$ and the associated standard deviation $\sigma_{X}$. We can notice that even under steady decarbonization, the ratio of cumulative emissions to cumulative GDP will remain higher than the annual carbon intensity. We also show in SI that larger fluctuations in the economy (larger $\sigma_{X}$) lead in average to a larger decoupling between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ because of phases of economic upturns and high carbon intensities followed by phases of recessions and low carbon intensities. In this case, reducing the annual carbon intensity has little effect on the path-dependent carbon intensity, yielding a larger degree of decoupling.
The trends for $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ are the same across all countries, with varying carbon intensities. For small GDP per capita, the relationship is almost linear and tends to a power law with an exponent slightly smaller than one at larger GDP per capita, as shown in Fig.2C. The lower this exponent, the more significant the efforts to decarbonize the economy.
Figure 2: A Path-dependent intensities $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ starting from 2000 and yearly carbon intensities $I(\tau)$ (consumption-based) for different countries. Data for path-dependent intensities and yearly carbon intensities are shown in dotted lines, theory for these quantities for which $C_{\tau}$ was computed from our model, and $I_{\tau}$ was obtained from data and is shown as a solid line (average), with the colored area representing standard deviation. Even if yearly carbon intensities tend to decrease (the carbon cost of $1USD$ invested decreases with time), path-dependent intensities decrease more slowly than intensities. Therefore, path-dependent intensities can almost be considered as constant on a yearly timescale. In practice, we use data on carbon intensities and our model, which incorporates country-specific measured values of the intrinsic growth rate, interest rate, and leverage, to estimate $C_{\tau}$. We report the mean average errors (MAE) and coefficients of determination between theory and data (path-dependent intensities) for all countries in the Methods section. B Comparing cumulative $CO_{2}$ emissions (data in dotted lines) for different countries (US, China, Denmark, and France) to the results of our stochastic model (in full lines, with standard deviation represented by the colored area). We report the mean average errors (MAE) and coefficients of determination between the theory and data (cumulative emissions per capita since 1998) for all countries in the Methods section. C Relationship between cumulative emissions per capita and cumulative GDP per capita from the data: we observe a clear increasing trend between both, with a slope that varies by country. Here, too, it is clear that China has the most carbon-intensive economy. Yet the US is responsible for the highest cumulative emissions, owing to its high cumulative GDP per capita. D Correlation between cumulative GDP and cumulative debt (left axis), and correlation between cumulative emissions and cumulative debt (right) axis. Cumulative GDP (circles) is mostly linear as a function of cumulative debt, except for Denmark, where the relationship between cumulative GDP and cumulative debt is slightly faster than linear. Concerning cumulative emissions (pentagons), the trend is initially linearly correlated with cumulative debt, then the correlation becomes slower than linear (in particular for China). Again, this is an effect of the decrease in carbon intensity with time.
By shifting consumption forward and increasing annual GDP per capita, debt increases cumulative emissions. This result is robust even if debt is used to decrease the annual carbon intensity, $I(\tau)$, because cumulative emissions are path-dependent.
Energetic and dissipative cost of an economy
Figure 3: A Effect of leverage: predictions for cumulative emissions (starting from 1998) depending on leverage in different countries. For each country, the color gradient indicates increasing leverage from $1.00$ to $1.05$, resulting in higher cumulative emissions per capita. A linearly decreasing carbon intensity has been assumed, with different parameters for each country [38], based on consumption-based emissions from 1998 to 2022. We also assume that the distributions of $\gamma$ remain constant for each country. We use $W=\left<\ln(\gamma_{\tau})\right$> and $\sigma_{X}$ (see values in Methods), as well as a fixed value of leverage, as inputs for the model. As expected, the larger the leverage, the greater the potential for GDP growth, leading to higher cumulative emissions despite decreasing carbon intensities. Changes by a few percent in leverage are enough to trigger changes of almost 11 order of magnitude in cumulative emissions per capita after $25$ years. B Effect of policies: Predictions for cumulative emissions (starting from 1998) depending on how strong the effort is to reduce carbon intensity. For each country, the color gradient corresponds to increasing $\left|dI_{\tau}/d\tau\right|$ (where $dI_{\tau}/d\tau$ is negative), indicating an increasing effort to reduce the carbon intensity of the economy. $\left|dI_{\tau}/d\tau\right|$ varies between 0 and $1.5$ times the current trend for each country. We observe that for high enough values of $\left|dI_{\tau}/d\tau\right|$, $I_{\tau}$ becomes negative (carbon capture), leading to decreasing cumulative emissions. The leverage is taken equal to the average leverage measured over the period 199820221998-2022 for each country.
Considering a period $t$, we can write the cumulative emissions $\mathcal{E}(t)$ as:
$\begin{split}\mathcal{E}(t)&=\sum_{\tau\leq t}\epsilon(C_{\tau},\tau),\end{split}$ (6)
assuming in a first approximation that yearly carbon intensity $I(\tau)$ does not depend explicitly on $C_{\tau}$, it has historically been measured as almost linear in most countries [38]. Therefore, writing $I(\tau)=I_{0}-\eta(\tau-\tau_{0})$,
$\mathcal{E}(t)=I_{0}\sum_{\tau\leq t}C_{\tau}-\eta\sum_{\tau\leq t}\tau C_{\tau}.$ (7)
Using this formula for the US $GDP$ data between 1990 and 2020, we obtain cumulative carbon emissions of $\mathcal{E}=7\times 10^{5}kg\,CO_{2}/\text{capita}$, close enough to the real value of $6\times 10^{5}kg\,CO_{2}/\text{capita}$ on Fig.2B.
We have analyzed data for the US deficit [42] and found that the added deficit in 2023 is $1.7\times 10^{12}USD$, resulting in a deficit per capita of $5.1\times 10^{3}USD$ in 2023. From this, we can obtain a value of the leverage in 2023 of $L=1.06$. If we use strategy B to describe the growth of the US economy, the intrinsic growth rate (that is, the growth rate minus debt-funded spending) is:
$\gamma_{eff,\tau}=\frac{C_{\tau+1}-C_{\tau}}{C_{\tau}}-(L-1).$ (8)
Studying this quantity allows for the decorrelation of the value added from credit to the added value (accounting for depreciation and production). On average, since 1980, this value is close to 0. Using data [43, 44, 45] for US GDP per capita since 1960, we can compute its growth rate. However, as mentioned earlier, we must distinguish the intrinsic growth rate from growth due to credits.
While the public deficit can stimulate economic activity within a Keynesian framework [46, 47], it also increases cumulative carbon emissions in proportion to the path-dependent carbon intensity. This highlights that even growth-stimulating fiscal policies contribute to emissions lock-in unless accompanied by structural reductions in carbon intensity, and raises the need to question growth targets themselves.
Usually, the goals to limit global warming are set in terms of total cumulative $CO_{2}$ emissions $\mathcal{E}$ [48, 38, 49, 50, 51], where a given scenario corresponds to a maximal amount of $CO_{2}$ emitted over a time domain. We can compare the data on cumulative $CO_{2}$ emissions with the model’s expected emissions and obtain a satisfactory match, as shown in Fig.1B. We can then use the model to predict emissions for the following years, as shown in Figs. 3A and B, depending on the leverage (used to maintain GDP growth) or growth rate $W$. We observe that cumulative emissions increase with both leverage and growth rate. Using the temperature anomaly as proportional to $CO_{2}$ emissions [48, 38], we can deduce that the temperature would rise accordingly. However, our results depend on the specific economies of the countries under study and do not capture cumulative emissions on a global scale, nor do they account for global changes in consumption habits.
Study of the solvency probability — the likelihood that debt can be paid back before capital depletion
We can relate the probability of bankruptcy to the probability that a biased random walk is smaller than a threshold at payback time $t_{p}$ [52, 53, 25] (the jumps do not have a zero mean, and the random walk is thus biased). We emphasize that bankruptcy does not correspond to the existence of a time where the debt is larger than the gross capital (or first passage time), but to the fact that at the time of payback $t_{p}$, the debt is larger than the gross capital.
Strategy A
We find a probability of solvency.
$P_{S}(t_{p})=\frac{1}{2}\left(1+\text{erf}\left(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}}\right)\right),$ (9)
where $m=\mathbb{E}(\ln(\gamma/\rho))=W-\ln(\rho)$. We plot this probability on Fig.2A, in the limit of large leverage (associated with higher risk). Interestingly, the result will strongly depend on the sign of $m$:
  • If $m>0$, $P_{S}(t_{p}\to\infty)\to 1$, faster as $\sigma_{X}$ decreases.
  • If $m<0$, $P_{S}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. In this case, the best course of action is to pay back as soon as possible.
  • If $m=0$, $P_{S}(t_{p}\to\infty)=\frac{1}{2}$.
Strategy B
Using strategy B, the solvency probability cannot be approximated as for strategy A. We obtain instead
$\begin{split}P_{S}(t_{p})=\mathbb{P}\left(X_{t_{p}}\geq 1-\frac{1}{L}\right),\end{split}$ (10)
with $X_{t_{p}}=\Pi_{t_{p}}-(L-1)\sum_{t\leq t_{p}-1}\Pi_{t}/\rho$ where $\Pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)/\rho$. We observe in Fig.4D that it is a decreasing function of time, as shown in the SI. Interestingly, this result holds true whenever $\left<\gamma\right><\rho$ and for low enough leverages. Thus, we recover the inequality that sets the stability of public debt with respect to GDP [34]. From this, we conclude that debt tends to rise faster on average than capital as long as the intrinsic growth rate is smaller than the interest rate $\rho$.
Typically, solvency is lower under strategy B because a higher risk must be taken to sustain a higher growth rate. For $L\sim 1$ we understand why this solvency probability (strategy B) is smaller than for strategy A, as the lower bound for solvency is larger. We also show in the SI that increasing leverage (thus relying more on credit) can be considered ”optimal” because it increases the growth rate of $C_{\tau}$, whereas the survival probability plateaus at large leverage. However, relying on such a strategy implies stronger carbon emissions, particularly for carbon intensive economies, as discussed above.
Effect of leverage on solvency (strategy B)
The risk associated with growth stochasticity can be evaluated using the probability of solvency. In particular, we can examine the impact of leverage and payback time on the probability of solvency. On Fig.4 we see that solvency probability generally decreases with both leverage and payback time in the range of leverage corresponding to real data. This means that debt is used as a means to increase short-term GDP, even if it results in increased long-term risk, which is measured by cumulative carbon emissions (depending on the economy’s carbon intensity) and a debt that grows faster than GDP.
In SI 2.D and 2.E, we further explore solvency in both models, and compute the optimal payback time for the post-payback growth rate conditional on solvency.
Figure 4: Illustration of trade-off between solvency and payback time. Solvency probability $P_{S}$ is shown for model B with the color scale as function of payback time $t_{p}$ and interest rate $\rho$. Solvency becomes less likely with time in this model whenever $\left<\gamma\right><\rho$ for all leverages $L$, i.e. when debt increases faster than the capital. This means that this strategy is not viable on long timescales, and is just a way to increase short-term GDP (resulting in increased carbon emissions).
Discussion
Our analysis reveals that debt-financed growth acts as a structural amplifier of emissions: leveraging $L$ boosts GDP $C_{\tau}$ and, given the slow evolution of carbon intensity $I(\tau)$, results in cumulative emissions that remain nearly proportional to cumulative GDP, confirming previous studies [37][54][55]. By separating the growth due to deficits from the intrinsic growth, we have confirmed that the economic growth of several countries is closely tied to deficits.
A second important observation is that although the intensity of carbon $I(\tau)$ generally declines due to technological improvements and efforts to develop clean energy, increasing the leverage $L$ always amplifies GDP $(C_{\tau})$, compensating technological gains in clean energy production. As a result, we predict that total emissions $\epsilon(t)$ will continue to increase in the coming decades.
We calibrated the model over multi-decade windows using observed macroeconomic data, extracting each country’s effective leverage $L$ (the average ratio of deficit-financed spending to GDP), the intrinsic growth rate $\gamma$ (the underlying growth rate absent new borrowing), and an interest rate $\rho$. This data-driven approach is not a free fit, but rather implements recorded deficits, public debt, and GDP to set model parameters. Therefore, the model’s predictions can be directly compared to the actual outcomes.
To our knowledge, this is the first model to explicitly link fiscal leverage with cumulative CO2 emissions and to calibrate it using multi-decade country-level data. Indeed, economic growth is typically treated as exogenous or as the result of a production function in Integrated Assessment Models (IAMs), such as DICE, REMIND, or MESSAGEix, without explicitly modelling credit dynamics [56, 57]. Similarly, the majority of climate-finance studies focus on transition risks or stranded assets [17], rather than quantifying the structural amplification of GDP growth and emissions by leverage. Our method is similar to macro-energy analyses that connect debt and energy consumption [2, 58]. However, it goes further by calibrating leverage $L$, intrinsic growth $\gamma$, and interest rate $\rho$ based on historical data. This allows for the direct prediction and validation of cumulative emissions trajectories. For all four considered countries, the model accurately tracks the cumulative CO2 emissions per capita alongside economic growth (Fig. 2B and SI), confirming that cumulative emissions are closely coupled with cumulative GDP, thus resulting in a persistent carbon lock-in effect. We can observe that, despite differences in decarbonization efforts, each additional unit of GDP has historically carried a largely invariant carbon cost over the long run (within the calibration period), carrying near-proportional relationships between total economic output and total emissions. For example, using a time-declining carbon intensity $I(t)$ for the US, the integration of our model from 1990 to 2020 yields cumulative per capita CO2 emissions of the order $7\times 10^{5}$ kg, in good agreement with the observed value of $\sim 6\times 10^{5}$ kg. A similar close agreement is found for China, France, and Denmark (Fig. 2B), as evidenced by the overlapping model trajectories and historical data points. This level of precision indicates that our leveraged growth model captures a first-order causal driver of emissions: debt-fuelled economic expansion. Hence, it can provide a partial explanation for why emissions have continued to increase in line with commodity production, despite gains in energy efficiency in many sectors. The differences between countries in economic structure and debt dynamics are reflected in the calibrated parameters and outcomes. We find that the United States is heavily dependent on deficit spending to sustain its growth. Indeed, since 1980, its intrinsic growth rate has averaged to only one percent per year, with substantial variability. In other words, most of the US GDP growth since 1980 can be attributed to credit expansion (public deficit and debt accrual) rather than to productivity gains [59]. Our calibration leads to an average leverage slightly above 1 for the US, rising to $L\approx 1.06$ in 2023 amid large federal spending. This credit-driven growth is reflected in the model’s ability in reproducing the persistent rise in cumulative U.S. CO2 emissions at the end of the 90s until just before the global financial crisis of 2008 (GFC) [60].
Quantitatively, the correspondence between modeled and observed cumulative CO2 is such that a simple linear regression of one against the other would yield a slope near 1, i.e. cumulative emissions are almost a linear function of cumulative GDP (see Fig.2C). The scatter of points around that line is small, thus reflecting primarily year-to-year economic variability rather than structural error in the model. This accordance is relevant, given that we did not calibrate the model by fitting it to emissions data, but rather derived emissions from independently calibrated economic parameters and an exogenous decline in carbon intensity. The observed results reveal that the coupling between economic output and emissions (illustrated by the Kaya identity, [61]) is grounded in the macro-financial structural dynamics of growth and investment.
An additional insight from our framework is that debt-driven growth imposes financial risks and constraints. Our stochastic analysis of the solvency probability (Fig. 4 and SI) shows that when the intrinsic growth rate $\langle\gamma\rangle$ falls below the interest rate $\rho$, the probability of long-term solvency decays towards zero, regardless of leverage $L$. In other words, maintaining growth through excessive borrowing ultimately becomes unsustainable and carries an endogenous risk of default. This result provides a natural constraint on fiscal stimulus strategies: even before critical climate transitions are reached, macro-financial dynamics can force deleveraging and major systemic destabilisations. From a control theory perspective, this identifies an optimal leverage level on the Pareto front, beyond which additional credit raises default risk more than it supports sustainable growth [62, 63, 23].
In this sense, the model provides a quantitative link between carbon debt and financial debt: high-leverage, high-emission growth paths not only increase cumulative CO2 emissions but also increase the risk of abrupt macroeconomic adjustments if interest rates rise or economic growth slows. Policies that are based on debt-financed growth should be assessed for their climate implications and the risk they pose of triggering financial instability and default.
One key limitation of the model is its tight coupling of public debt and GDP. We assume that each country’s government borrows a fixed fraction of GDP each year (the leverage $L$ above 1 corresponds to a primary deficit ratio) and that this borrowing directly and immediately adds to that year’s GDP. In reality, the relationship between debt and growth is more complex: private credit creation (by banks and firms) can also drive variability, and not all deficit spending translates into GDP in the short run (in fact, the fiscal multiplier can vary). Our model effectively treats deficit-financed spending as equally productive as other spending, whereas inefficient investments or financial sector leakages could reduce the growth payoff of debt. In addition, identifying the primary deficit with new capital investment is an oversimplification. In fact, governments may borrow for many reasons, and not all borrowing fuels productive capacity. A more advanced model would require integrating both public and private debt dynamics and the ability to distinguish between productive investment and pure demand stimulus.
Another important limitation is the model’s lack of sectoral differentiation in its treatment of emissions. We model each country as a single homogeneous production unit, with a single aggregate carbon intensity of output, in a manner similar to the Nordhaus model [64, 65, 15, 66]. This implies that we do not account for changes in the composition of GDP from high-carbon sectors (such as heavy industry) to low-carbon sectors (such as services). For example, within a country, some industries can rapidly decarbonize (such as renewables in power generation), while others lag behind (such as air travel or defense) [67, 68]. Our one-sector approach averages these dynamics out. Additionally, we assume that the same carbon intensity function $I(t)$ applies regardless of the amount produced, thereby ignoring any nonlinear feedback (such as resource constraints, climate risks, or price effects) that may occur. In reality, a large increase in credit use (as in China’s case) can temporarily increase carbon intensity by increasing investment in construction and infrastructure. Our model partially reflects this by calibrating $I(t)$ piecewise (using historical data to set $I_{0}$ and a decline rate $\eta$). China’s financial system structure may partially explain the apparent decoupling between GDP growth and territorial CO2 emissions. Unlike market-based systems, in which credit allocation is determined by profit expectations, China’s state-dominated banking sector sets lending priorities based on policy objectives. A large-scale financing of renewables, electrified transport, and grid infrastructure is facilitated by the government’s determination of which sectors are eligible for credit. According to the model, this centralised allocation temporarily reduces the leverage–emissions gap by increasing the effective decarbonization rate $\beta$ in comparison to the growth rate $\gamma$.
Denmark presents another informative and distinctive boundary case for our framework, owing to its limited public borrowing and persistent fiscal surpluses. As illustrated in the SI, Denmark’s territorial emissions have decreased, while its consumption-based emissions have remained relatively constant, indicating that no actual decoupling has occurred. This discrepancy suggests that fiscal restraint alone is insufficient to guarantee a reduced carbon footprint when emissions are outsourced through trade. In the SI, we compare territorial and consumption-based estimates (Fig. S4) and find that offshoring of emissions can significantly alter apparent decoupling trends (a phenomenon also referred to as carbon leakage [69]). Despite their high consumption-based footprint, countries that relocate carbon-intensive production abroad appear to decrease their emissions-to-GDP ratio [70]. In a similar spirit, network-based allocation methods developed by Van den Ende et al. (2025) offer a rigorous approach to quantify responsibility for emissions along global value chains, which complements territorial accounting [71].
It’s worth noticing that our analysis does not currently include private credit and investment flows, which can be significant drivers of emissions and capital formation [72]. The current representation of the leverage term $L_{t}$ is based on public fiscal deficits, which account only for the government’s contribution to output growth through borrowing. Nevertheless, alternative data sources, such as total credit to the non-financial sector or capital market issuance, could be used to illustrate the intensification of growth through private investment flows. This expansion would enable the model to quantitatively describe mechanisms, such as those specifically identified in recent analyses on institutional investors and insurers, that persist in financing the expansion of fossil fuels despite climate commitments [73][74].
Financial debt and carbon debt are two sides of the same anticipatory system: the former brings future wealth forward, while the latter consumes the energy of the future. Kelly’s model, when scaled up to the macroeconomic level, shows that there are both physical and financial limits to debt-driven growth. It is an important message because the idea that there are limits to growth, even when allowing for innovation, is not widely accepted among economists [75]. Here, we show that as long as credit rests on the promise of material expansion, each euro borrowed creates a double debt — a financial and a climatic one. The great challenge of the twenty-first century will be to decouple prosperity from the credit lever, so that the growth of well-being is no longer tied to the growth of carbon.
Methods
We report the parameters used in the model to simulate the evolution of the $GDP$ in Table 1. We averaged their values over the interval 199820221998-2022 to use as inputs in our model.
China US Denmark France
$\left<L_{\tau}\right$> $1.025\pm 0.023$ $1.051\pm 0.040$ $0.997\pm 0.029$ $1.040\pm 0.018$
$\left<\rho_{\tau}\right$> $1.131\pm 0.101$ $1.013\pm 0.051$ $1.018\pm 0.237$ $1.008\pm 0.088$
$W$ $088\pm 0.086 0.0$$18\pm 0.059 0.0$$32\pm 0.089 0.0$$020\pm 0.090$
$\left<\gamma_{\tau}\right$> $1.096\pm 0.095$ $0.983\pm 0.056$ $1.036\pm 0.091$ $0.983\pm 0.088$
Table 1: Values of the different parameters.
We also report the coefficients of determination $R^{2}$ and rescaled Mean Absolute Errors obtained between data and our model for Fig.2 A and B in table.2 and table.3. The model is used to simulate the evolution of the GDP with its related stochasticity, and uses measured yearly carbon intensities as an input to simulate path-dependent intensities $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$.
China US Denmark France
$R^{2}_{\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)}$ $0.985$ $0.938$ $0.981$ $0.982$
$R^{2}_{\mathcal{E}(\tau)}$ $0.980$ $0.996$ $0.982$ $0.977$
Table 2: Coefficients of determination $R^{2}$ between data and theory for path-dependent intensities and cumulative emissions per capita, for different countries. With $(X_{n})$ being the real data and $X_{n}^{\text{simu}}$ the simulated data, we have $R^{2}_{X}=1-\sum\sum_{n=1}^{N_{m}}\left((X_{n}-X_{n}^{\text{simu}})/(X_{n}-\left<X\right>)\right)^{2}$.
China US Denmark France
$MAE_{\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)}/\left<\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\right$> $0.032$ $0.035$ $0.027$ $0.017$
$MAE_{\mathcal{E}(\tau)}/\left<\mathcal{E}(\tau)\right$> $0.092$ $0.027$ $0.057$ $0.057$
Table 3: Rescaled Mean Absolute Error MAEMAE between data and theory for path-dependent intensities and cumulative emissions per capita, for different countries. This quantity is computed as $MAE_{X}/\left<X\right>=\left(\sum_{n=1}^{N_{m}}|X_{n}-X_{n}^{\text{simu}}|\right)/\left(\sum_{n=1}^{N_{m}}X_{n}\right)$ with $N_{m}$ the number of measurements.
Acknowledgments
We are grateful to Eric Galbraith, Julia Steinberger, Nico Wunderling, Jeff Althouse, and André de Palma for their insightful feedback and discussions, which greatly strengthened this work. S.M. acknowledges support from the 2025 FIR–PRI (Best Research Grant) for their PhD project.
Appendix A Gambling model and its connection to macroeconomic models for public debt
In this section, we present the gambling model studied in this paper, which is an extension of Kelly’s gambling model, and then, we explain how this extended Kelly’s model relates to known macroeconomic models for public debt.
A.1 Definition of the gambling model
We assume that at each time step $\tau$ (we use discrete time), a gambler can borrow an amount $B_{\tau}$ with interest rate $\rho_{\tau}$. The interest rate $\rho_{\tau}$ is the factor by which the amount borrowed is multiplied at each time step, in other words, the debt due to the amount borrowed at time $\tau^{\prime}$, is multiplied by $\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}$ if it has not been paid back at time $\tau$. In general, the interest rate is larger than 11 ($\rho>1$) so that the debt keeps increasing, and the borrower ends up paying more than the amount that was borrowed. Then the gross capital $C_{\tau+1}$, and debt $D_{\tau+1}$ after one step are (until bankruptcy occurs if $C_{\tau+1}\leq 0$):
$C_{0}$ $=B_{0}+C_{\text{ini}}=LC_{\text{ini}}$ (11)
$C_{\tau+1}$ $=\gamma_{\tau}C_{\tau}+B_{\tau+1}-\left(\sum_{\tau^{\prime}=0}^{\tau}\delta_{\tau,t_{p}(\tau^{\prime})}\right)D_{\tau+1}$ (12)
$D_{0}$ $=B_{0}=(L-1)C_{\text{ini}}$ (13)
$D_{\tau}$ $=\sum_{\tau^{\prime}=0}^{\tau-1}\Theta(t_{p}(\tau^{\prime})-\tau)\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}B_{\tau^{\prime}},$ (14)
where $\gamma_{\tau}$ is stochastic and depends on time (in a classical Kelly’s model, where $x$ represents the winner of time $\tau$, happening with a probability given by the probability vector $\boldsymbol{p}$, $\boldsymbol{b}$ is the gambling strategy and $\boldsymbol{r}$ is the inverse-odds vector, we would get that $\gamma_{\tau}=b_{x,\tau}/r_{x,\tau}$ which is a random variable which distribution depends on $\boldsymbol{p}$).. The first term in the first equation represents the standard evolution in Kelly’s model, namely the contribution of gambling at time $\tau$: if $\gamma_{\tau}>1$, the gross capital increases provided that there is no borrowing or payback at this time. The second term represents the amount borrowed at time $\tau$; it can be 0 if the borrower does not borrow at this step, or a positive amount. The third term represents the possible payback of the debt (debt clearing) at time $\tau+1$, and will either be equal to 0 (no payback), or to $D_{\tau+1}$. $C_{\text{ini}}$ is the amount invested initially by the gambler, and $B_{0}$ the initial amount borrowed. The leverage is defined as $L=(B_{0}+C_{\text{ini}})/C-i$ and quantifies the dependency of the gambler on an initial borrowing from the bank.
In addition $t_{p}(\tau^{\prime})$ indicates the time when the amount borrowed at $\tau^{\prime}$ is payed back, and the Heaviside function $\Theta$ function allows to know if the payback has occurred at time $\tau$ (indeed $\Theta(t_{p}(\tau^{\prime})-\tau)=0$ if $\tau>t_{p}$, meaning that the debt is erased after it has been payed back). If a payback happens at time $\tau$, it means that $\exists\tau^{\prime},\tau=t_{p}(\tau^{\prime})$, and thus $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=1$, otherwise $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=0$.
As explained above, without debt, the gross capital cannot hit 0 in Kelly’s model, and bankruptcy never occurs. In our updated view of this model, the capital can become negative if the debt exceeds the capital when the debt is repaid. Therefore solvency means that the debt is always paid back at a time where the gross capital is larger ($C_{\tau}>D_{\tau+}$). We call $C_{\tau}-D_{\tau}$ the net capital, whose sign indicates if bankruptcy will happen in case of payback.
A.2 Single loan model (strategy A)
For strategy A, we assume that the gambler initially borrows money ($B_{0}>0$) but never borrows afterwards, $\forall\tau\geq 1,B_{\tau}=0$. This strategy is more like that of an individual taking out a loan to buy a house. In practice, the gambler can borrow $B_{0}$ and invest $C_{\text{ini}}$ from its personal capital, and a key quantity will be the leverage $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model with an initial investment of $B_{0}+C_{\text{ini}}$, and the debt evolves with an interest rate $\rho$
$\begin{split}C_{\tau}&=(B_{0}+C_{\text{ini}})\prod_{1\leq\tau^{\prime}\leq\tau}\gamma_{\tau^{\prime}}=C_{\text{ini}}L\prod_{1\leq\tau^{\prime}\leq\tau}\gamma_{\tau^{\prime}}.\\ D_{\tau}&=B_{0}\rho^{\tau}=C_{\text{ini}}(L-1)\rho^{\tau}\end{split}$ (15)
Then at the time of payback, for $\tau=t_{p}$, we find:
$\begin{split}C_{t_{p}}&=B_{0}\left(\prod_{1\leq\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}-\rho^{t_{p}}\right)+C_{\text{ini}}\prod_{\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}\\ &=C_{\text{ini}}\left[L\left(\prod_{1\leq\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}-\rho^{t_{p}}\right)+\rho^{t_{p}}\right]\end{split}$ (16)
Obviously, this amount can be either negative or positive, which could not happen within the purely multiplicative Kelly’s model. Bankruptcy occurs if $C_{t_{p}}$ is negative or equivalently
$\prod_{\tau^{\prime}\leq t_{p}}\frac{\gamma_{\tau^{\prime}}}{\rho}<1-\frac{1}{L},$ (17)
meaning that the larger $L$ is, the more bankruptcy is likely to occur. In average over the joint probability $p(x_{1},...,x_{t_{p}})$, using the fact that the races are independent, we find:
$\begin{split}\left<C_{t_{p}}\right>&=C_{\text{ini}}\left[L\left(\sum_{x_{1},...,x_{t_{p}}}\mathbb{P}(x_{1},...x_{t_{p}})\prod_{i=1}^{t_{p}}\gamma_{x_{i}}-\rho^{t_{p}}\right)+\rho^{t_{p}}\right]\\ &=C_{\text{ini}}\left[L\left(\sum_{x_{1},...,x_{t_{p}}}\prod_{i=1}^{t_{p}}\mathbb{P}(x_{i})\gamma_{x_{i}}-\rho^{t_{p}}\right)+\rho^{t_{p}}\right]\\ &=C_{\text{ini}}\left[L\left(\prod_{i=1}^{t_{p}}\sum_{x_{i}}\mathbb{P}(x_{i})\gamma_{x_{i}}-\rho^{t_{p}}\right)+\rho^{t_{p}}\right]\\ &=C_{\text{ini}}\left[L\left(E^{t_{p}}-\rho^{t_{p}}\right)+\rho^{t_{p}}\right],\end{split}$ (18)
where $E=\mathbb{E}(\gamma_{x})$ is the average earning after each race (in the following we will consider the case $E>1$). We can also define the average short-term growth rate
$w(t_{p})=\frac{1}{t_{p}}\ln\left(\frac{\left<C_{t_{p}}\right>}{LC_{\text{ini}}}\right)=\frac{1}{t_{p}}\ln\left(E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}\right),$ (19)
which is the average growth rate of capital relative to the initial capital $LC_{\text{ini}}$ just after the debt is repaid, assuming no bankruptcy. This way we can write that $C_{t_{p}}=LC_{\text{ini}}\exp(w(t_{p})t_{p})$. The larger this quantity is, the faster the capital will grow after payback. If the gambler is thinking in short terms, this is a quantity to maximize. Interestingly we can bound this quantity (because $(1-1/L)\rho^{t_{p}}>0$) :
$\forall t_{p},\ln(E)\geq w(t_{p})$ (20)
If bankruptcy does occur, the game stops. Otherwise, the gambler can keep playing, and for $\tau>t_{p}$:
$C_{\tau}=C_{\text{ini}}\left[L\left(\prod_{\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}-\rho^{t_{p}}\right)+\rho^{t_{p}}\right]\prod_{t_{p}<\tau^{\prime}\leq\tau}\gamma_{\tau^{\prime}}.$ (21)
From this we see that the gross capital is well defined for $(1/\rho^{t_{p}})\prod_{\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}>1-1/L$. Under this assumption (no bankruptcy), the growth rate conditioned on solvency is defined as:
$\begin{split}W&=\lim_{\tau\to\infty}\frac{\ln(C_{\tau})}{\tau}=\lim_{\tau\to\infty}\frac{1}{\tau}\left(\ln\left(\prod_{t_{p}<\tau^{\prime}\leq\tau}\gamma_{\tau^{\prime}}\right)+\ln\left(C_{\text{ini}}\left[L\left(\prod_{\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}-\rho^{t_{p}}\right)+\rho^{t_{p}}\right]\right)\right)\\ &=\lim_{\tau\to\infty}\frac{1}{\tau}\left(\ln\left(\prod_{t_{p}<\tau^{\prime}\leq\tau}\gamma_{\tau^{\prime}}\right)\right)\\ &=\mathbb{E}\left({\ln(\gamma_{x})}\right),\end{split}$ (22)
where the equivalence results from the weak law of large numbers, and the fact that the second term is a constant with respect to the time $\tau$ (thus disappears when multiplied by $1/\tau$ for large times). Note that due to the concavity of $\ln$, we always have (independent of the probability distribution $p_{x}$):
$\ln(E)\geq W$ (23)
However, there is a non-zero probability of hitting bankruptcy when the debt is paid back (at $t_{p}$). The probability of hitting bankruptcy is:
$P_{B}(t_{p})=\mathbb{P}\left(\prod_{\tau^{\prime}\leq t_{p}}\frac{\gamma_{\tau^{\prime}}}{\rho}\leq 1-\frac{1}{L}\right)=\mathbb{P}\left(\sum_{\tau^{\prime}\leq t_{p}}\ln(\frac{\gamma_{\tau^{\prime}}}{\rho})\leq\ln\left(1-\frac{1}{L}\right)\right).$ (24)
Within Kelly’s model, and with a strategy corresponding to Kelly’s optimum $b_{x}=p_{x}$, we obtain stochastic evolutions of the gross capital on Fig.1 from the main text. We observe that some trajectories hit bankruptcy at the time of payback, while others survive and keep increasing, depending on the history of results for $\tau\leq t_{p}$.
If we fix the probability distribution $p_{x}=[0.03,0.20,0.39,0.12,0.26]$ and the inverse odds $r_{x}=[0.15,0.39,0.04,0.21,0.21]$ using Kelly’s optimal strategy $b_{x}=p_{x}$, we plot the simulated solvency probability for different interest rates on Fig.10 in the limit of large leverage $L=15$. Depending on the value of the interest rate, the solvency probability can either increase (low interest rates) or decrease (large interest rates) with the payback time $t_{p}$.
A.3 Periodic borrowing model (strategy B)
In this section, we assume that the gambler borrows money at each time step, with a leverage $L_{\tau}$ (depending on time a priori), which we refer to as strategy B. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model, with a modified stochastic growth rate, and the debt still evolves with an interest rate $\rho$. The gambler might want to partially pay back the debt at each step, but we ignore this possibility for now. Before the debt is paid, we have the equations modified in the form
$\begin{split}C_{\tau+1}&=\left(\gamma_{\tau}+(L_{\tau}-1)\right)C_{\tau}\\ D_{\tau+1}&=\rho D_{\tau}+(L_{\tau}-1)C_{\tau},\end{split}$ (25)
where the second term in the first equation comes from credit, as well as the second term in the second equation. $G_{t}$ corresponds to government expenditures, while $T_{t}$ corresponds to taxation.
We assume that the leverage is constant in time in the following, and justify this assumption on Fig.5. Therefore, if we introduce $\tilde{D}_{\tau}=D_{\tau}/\rho^{\tau}$
$\begin{split}C_{\tau}&=C_{\text{ini}}L\prod_{1\leq\tau^{\prime}\leq\tau}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\\ \tilde{D}_{\tau+1}&=\tilde{D}_{\tau}+\frac{L-1}{\rho^{\tau+1}}C_{\tau},\end{split}$ (26)
and thus, from the second equation
$\tilde{D}_{\tau}=\tilde{D}_{0}+\sum_{\tau^{\prime}=0}^{\tau-1}\frac{L-1}{\rho^{\tau+1}}C_{\tau}.$ (27)
And finally, we find an expression for the debt
$D_{\tau}=\rho^{\tau}\left(D_{0}+\sum_{\tau^{\prime}=0}^{\tau-1}\frac{L-1}{\rho^{\tau+1}}C_{\tau}\right).$ (28)
So that we obtain
$\begin{split}C_{\tau}&=C_{\text{ini}}L\prod_{1\leq\tau^{\prime}\leq\tau}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\\ D_{\tau}&=\rho^{\tau}\left(C_{\text{ini}}(L-1)+\sum_{0\leq t\leq\tau-1}\frac{L-1}{\rho^{t+1}}C_{t}\right)=C_{\text{ini}}(L-1)\rho^{\tau}\left(1+\sum_{0\leq t\leq\tau-1}\frac{L}{\rho^{t+1}}\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\right).\end{split}$ (29)
Then at the time of payback, for $\tau=t_{p}$, we find:
$\begin{split}C_{t_{p}}&=C_{\text{ini}}L\prod_{1\leq\tau^{\prime}\leq t_{p}}\left(\gamma_{\tau^{\prime}}+(L-1)\right)-C_{\text{ini}}(L-1)\rho^{t_{p}}\left(1+\sum_{0\leq t\leq t_{p}-1}\frac{L}{\rho^{t+1}}\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\right)\\ &=C_{\text{ini}}\rho^{t_{p}}\left[L\left(\prod_{1\leq\tau^{\prime}\leq t_{p}}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)-1-\sum_{0\leq t\leq t_{p}-1}\frac{L-1}{\rho}\prod_{1\leq\tau^{\prime}\leq t}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\right)+1\right]\end{split}$ (30)
Therefore, using the same method as previously, we obtain the average
$\left<C_{t_{p}}\right>=C_{\text{ini}}\rho^{t_{p}}\left[L\left(\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}-1-\frac{L-1}{\rho}\frac{1-\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}}{1-\frac{E+(L-1)}{\rho}}\right)+1\right]$ (31)
This sets the average short term growth rate in this case
$w(t_{p})=\frac{1}{t_{p}}\ln\left(\rho^{t_{p}}\left(\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}-1-\frac{L-1}{\rho}\frac{1-\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}}{1-\frac{E+(L-1)}{\rho}}+\frac{1}{L}\right)\right)$ (32)
This way we can also study the solvency probability for strategy B, which is
$\begin{split}P_{B}(t_{p})&=\mathbb{P}\left(\prod_{1\leq\tau^{\prime}\leq t_{p}}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)-\sum_{t\leq t_{p}-1}\frac{L-1}{\rho}\prod_{0\leq 1\leq\tau^{\prime}\leq t}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\leq 1-\frac{1}{L}\right).\end{split}$ (33)
A.4 Relation to public debt models
In this section, we explicitly relate the model for the public debt of [33] to the gambling model outlined above. For the public debt, one usually studies debt in terms of the debt-to-GDP ratio. We call $C_{\tau}$ the GDP at time $\tau$, which can be split in different sectors : consumption $U_{\tau}$, investments $I(\tau)$, government spending $G_{\tau}$, exports $X_{\tau}$ and imports $M_{\tau}$ (setting $X_{\tau}-M_{\tau}$ to be the net exports). The primary deficit represents the value added to the public debt (apart from interests), and is split as government spending minus taxation $T_{\tau}$.
$\begin{split}C_{\tau}&=U_{\tau}+I(\tau)+G_{\tau}+(X_{\tau}-M_{\tau})\\ D_{\tau+1}&=\rho D_{\tau}+G_{\tau+1}-T_{\tau+1},\end{split}$ (34)
To map this with the previous model requires assuming that the primary deficit $G_{\tau}-T_{\tau}$ is a fraction of the GDP, borrowed to increase the yearly GDP. Even if the mechanism is more complex, and debt usually motivates carbon-intensive activities based on the assumption that taxes will eventually generate surpluses for the government. If we want to identify with our previous model, we can write that
$C_{\tau+1}=\gamma_{\tau}C_{\tau}+G_{\tau+1}-T_{\tau+1}.$ (35)
This leads to the identity
$\gamma_{\tau}-1=\frac{\left(T_{\tau+1}+U_{\tau+1}+I(\tau+1)+X_{\tau+1}-M_{\tau+1}\right)-C_{\tau}}{C_{\tau}},$ (36)
which means that $\gamma_{\tau}$ encapsulates relative growth due to non-governmental activities, particularly if taxation is an increasing function of economic fluxes. With such definitions, we obtain
$\begin{split}C_{\tau+1}&=\gamma_{\tau}C_{\tau}+\left(G_{\tau+1}-T_{\tau+1}\right)\\ D_{\tau+1}&=\rho D_{\tau}+\left(G_{\tau+1}-T_{\tau+1}\right),\end{split}$ (37)
which we can identify with the previous model a priori, under the condition that we define a leverage
$L_{\tau}=1+\frac{G_{\tau+1}-T_{\tau+1}}{C_{\tau}},$ (38)
which thus corresponds to the comparison between the primary deficit of year $\tau+1$ to the GDP of year $\tau$. Indeed, we can introduce the deficit $d_{\tau+1}=G_{\tau+1}-T_{\tau+1}$, and we get $L_{\tau}=1+d_{\tau+1}/C_{\tau}$. To assess the validity of our hypothesis that this leverage is fixed for a given country, we show this value from data for the US, China, France, and Denmark in Fig.5. We observe a small variation (of a few percent) over the period of study, indeed.
Figure 5: Comparing primary deficit to GDP for different countries. In dotted lines we show the data, in full lines the average values and the colored areas represent standard deviations. We see that the typical variations are of a few percent. The trend for the US, China and France seems to be a slight increase in the ratio. We also notice that the ratio is negative on average for Denmark, meaning that the government typically has surpluses.
A.5 Relation to the Keen model
The Keen model [16] is a model that describes the dynamics of macroeconomic variables. In that model, the equations for $C_{\tau}$ and $D_{\tau}$, take the following form in terms of the total growth rate $g_{\tau}$, the total interest rate $r_{\tau}$, the investment to output ratio $\kappa_{\tau}$, the wage to labor ratio $\omega_{\tau}$, and the output to labor ratio $a_{\tau}$:
$\begin{split}C_{\tau+1}-C_{\tau}&=g_{\tau}C_{\tau}\\ D_{\tau+1}-D_{\tau}&=r_{\tau}D_{\tau}+\left(\kappa_{\tau}+\frac{\omega_{\tau}}{a_{\tau}}-1\right)C_{\tau}.\end{split}$ (39)
This leads to the following mapping to our model
$\begin{split}\gamma_{\tau}&\sim g_{\tau}-\left(\kappa_{\tau}+\frac{\omega_{\tau}}{a_{\tau}}\right)\\ \rho_{\tau}&\sim 1+r_{\tau}\\ L_{\tau}&\sim\left(\kappa_{\tau}+\frac{w_{\tau}}{a_{\tau}}\right).\end{split}$ (40)
In practice $\omega_{\tau}/a_{\tau}$ is the wage to output ratio, and $\kappa_{\tau}$ represents the share of output dedicated to investments. Following Fig.5, this suggests that the yearly variation of $\kappa_{\tau}+\omega_{\tau}/a_{\tau}$ is weak. The stochasticity of our model is thus encapsulated in $g_{\tau}$. This also gives insights on the variables, $L_{\tau}$ accounts for investments and labour.
Appendix B Results
B.1 Lock-in of carbon emissions
Assuming that an external bank lends money each year at the cost of increasing the debt, we derive the evolution of a country’s economy in terms of its GDP $C_{\tau}$. In practice, the capital will be used to invest in energetically costly activities. This means that we can associate to $C_{\tau}$ a function $e(C_{\tau})$ which represents the energetic cost of the GDP $C_{\tau}$. This should be an increasing function of the capital, and thus its average should increase with $L$ and $C_{\text{ini}}$. There exists a relationship between energy and economical value in dollars for different sectors, explored with Leontief’s input output analysis in [76]. If we also assume that emission, or dissipation $\epsilon$ is an increasing function of energy consumption, we ultimately have that $\epsilon$ is an increasing function of the capital $C_{\tau}$ (which is observed in practice [37]). A second order relationship [37] $\epsilon(C_{\tau})=\alpha_{0}+\alpha_{1}C_{\tau}+\alpha_{2}C_{\tau}$ can be used to model emissions encapsulate that emissions tend to increase faster as $GDP$ increases. The parameters $\alpha_{0,1,2}$ may depend on time, for instance if more renewable energies are used over time. We rather use the carbon intensity of an economy $I$ [38], representing the carbon cost of $1USD$ (which depends on a country’s activities, energy mix, …. This is given by Kaya’s relationship [38, 39, 14, 40] -
$\epsilon(\tau,C_{\tau})=C_{\tau}\times I(\tau).$ (41)
The cumulative emissions of a country depend on the history of its economy, we refer to this effect as the lock-in effect, and measure it using the path-dependent intensity
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)=\frac{\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})C_{\tau^{\prime}}}{\sum_{\tau^{\prime}=0}^{\tau}C_{\tau^{\prime}}},$ (42)
B.2 Energetic cost of debt
On Table.4 we report the different values extracted from data, for different countries, averaged over the period 199820221998-2022. To obtain $L,\rho,W$, we use the following equations and data for $GDP$, debt, and deficit
$\begin{split}L_{\tau}&=1+\frac{d_{\tau+1}}{C_{\tau}}\\ \rho_{\tau}&=\frac{D_{\tau+1}-d_{\tau+1}}{D_{\tau}}\\ \gamma_{\tau}&=\frac{C_{\tau+1}-d_{\tau+1}}{C_{\tau}},\end{split}$ (43)
where $d_{\tau+1}$ is the deficit of the state during year $\tau$ for year $\tau+1$, and the fact that $W=\left<\gamma_{\tau}\right$>.
China US Denmark France
$\left<L_{\tau}\right$> $1.025\pm 0.023$ $1.051\pm 0.040$ $0.997\pm 0.029$ $1.040\pm 0.018$
$\left<\rho_{\tau}\right$> $1.125\pm 0.023$ $1.010\pm 0.051$ $1.014\pm 0.155$ $1.002\pm 0.091$
$W$ $088\pm 0.086 0.0$$18\pm 0.059 0.0$$32\pm 0.089 0.0$$020\pm 0.090$
Table 4: Values of the different parameters.
Considering a period $t$ we can write the cumulated emissions $\mathcal{E}(t)$ and resource consumptions $\mathcal{R}(t)$ as
$\begin{split}\mathcal{E}(t)&=\sum_{\tau\leq t}\epsilon(C_{\tau},\tau)\\ \mathcal{R}(t)&=\sum_{\tau\leq t}r(C_{\tau},\tau),\end{split}$ (44)
where $\epsilon(C,\tau)$ are the emissions of investing $C$ at time $\tau$, and $r(C,\tau)$ is the total use in resource when investing $C$ at time $\tau$. For the US, we obtain Fig.7 with the coefficients of determination $R^{2}=0.82$ for GDP and $R^{2}=0.79$ for debt. To ensure the model is working, we also use it to model GDP and debt for other countries and estimate their associated emissions. For France, we get the coefficients of determination $R^{2}=0.53$ for GDP and $R^{2}=0.86$ for debt. For China, we get the coefficients of determination $R^{2}=0.91$ for GDP and $R^{2}=0.97$ for debt. For Denmark, we obtain coefficients of determination $R^{2}=0.62$ for GDP and a negative $R^{2}$ for debt, indicating that the model fails to capture the strong variations and changes in sign of leverage. We observe a greater discrepancy between predictions and data for debt in Denmark. This is because Denmark usually runs a surplus and partially repays the debt. Therefore, the predicted debt exceeds the actual debt, which usually decreases. We report the coefficients of determination and rescaled mean absolute errors on table.5.
China US Denmark France
$MAE_{GDP}/\left<{GDP}\right$> $0.155$ $0.081$ $0.116$ $0.119$
$MAE_{\text{Debt}}/\left<\text{Debt}\right$> $0.141$ $0.203$ $0.172$ $0.119$
$R^{2}_{GDP}$ $0.91$ $0.82$ $0.62$ $0.53$
$R^{2}_{\text{Debt}}$ $0.97$ $0.79$ 1.22-1.22 $0.86$
Table 5: Rescaled Mean Absolute Error MAEMAE and coefficient of determination $R^{2}$ between data and theory for GDP and debt, for different countries. $MAE_{X}/\left<X\right>=\left(\sum_{n=1}^{N_{m}}|X_{n}-X_{n}^{\text{simu}}|\right)/\left(\sum_{n=1}^{N_{m}}X_{n}\right)$ with $N_{m}$ the number of measurements and $R^{2}_{X}=\sum\sum_{n=1}^{N_{m}}\left((X_{n}-X_{n}^{\text{simu}})/(X_{n}-\left<X\right>)\right)^{2}$.
Another important point to study is the difference between consumption-based emissions, meaning that some countries may display a comparatively low carbon intensity by only accounting for territorial emissions. However, when accounting for exported emissions, the carbon intensity increases. This is illustrated with predictions using our model in Fig.6.
Figure 6: Comparing territorial emissions and consumption-based emissions using our model for predictions. Territorial emissions represent carbon emissions produced on the territory of a given country, while consumption-based counts emissions due to activities directly used by a given country. For France and Denmark, we observe that territorial emissions are lower than consumption-based. Predictions are performed by running the model with the distributions of values for leverage $L$, $W$, $\rho$, and carbon intensity over the period 1998-2022. Carbon intensities are assumed to be linear in time, with a constant decarbonation rate corresponding to that of the period 1998-2022 (which is different for territorial emissions and consumption-based emissions).
Figure 7: Comparison to data for GDPGDP per capita, and debt per capita between 2000 and $20$ for all countries. This is not a fit, but only using values for the deficit, GDP and debt to compute the parameters of the model.
We also realize a plot similar to that of Fig.3 in the main text, showing predictions of the model for various intrinsic growth rates on Fig.8.
Figure 8: Predictions for cumulative emissions depending on the long-term growth rate $W$ in different countries. We also used our stochastic model and assume a linearly decreasing carbon intensity (with different parameters for each country). We assume that the leverage remains the same in each country.
B.3 Comparing path dependent intensities to yearly intensities
If we want to compare $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ to $I(\tau)$ in general, we get
$\begin{split}\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)=\frac{\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau)\right)C_{\tau^{\prime}}}{\sum_{\tau^{\prime}=0}^{\tau}C_{\tau^{\prime}}},\end{split}$ (45)
from which we can easily see that if $\tau\to I(\tau)$ is a monotonically decreasing function, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ will always be larger than $I(\tau)$. $I(\tau)$ is a decreasing function in many countries due to changes in energy use, often referred to as innovation. We can then compute the evolution of the quantity , defining $S_{\tau}=\sum_{\tau^{\prime}=0}^{\tau}C_{\tau^{\prime}}$
$\begin{split}\mathfrak{I}_{\tau+1}\left(\{C_{\tau}\}\right)-\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)&=\frac{1}{S_{\tau}S_{\tau+1}}\left(\sum_{\tau^{\prime}=0}^{\tau+1}I(\tau^{\prime})S_{\tau}C_{\tau^{\prime}}-\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})S_{\tau+1}C_{\tau^{\prime}}\right)\\ &=\frac{1}{S_{\tau}S_{\tau+1}}\left(I(\tau+1)C_{\tau+1}S_{\tau}+\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})C_{\tau^{\prime}}\left(S_{\tau}-S_{\tau+1}\right)\right)\\ &=\frac{C_{\tau+1}}{S_{\tau+1}}\left(I(\tau+1)-\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\right).\end{split}$ (46)
The last equation leads to the conclusion that the path dependent intensity is decreasing with time if and only if $I(\tau+1)<\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ that is if and only if
$\sum_{\tau^{\prime}=0}^{\tau}\left(I_{\tau^{\prime}}-I(\tau+1)\right)C_{\tau^{\prime}}>0.$ (47)
This means that as long as the yearly carbon intensity is a decreasing function, so is the path dependent intensity. For a non monotonic intensity, the years with highest $C_{\tau}$ will prevail in the sum. From this we can also deduce the variations of $\Delta_{\tau}:=\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)$,
$\begin{split}\Delta_{\tau+1}-\Delta_{\tau}&=\frac{C_{\tau+1}}{S_{\tau+1}}\left(I(\tau+1)-\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\right)-(I(\tau+1)-I(\tau))\\ &=\frac{S_{\tau}}{S_{\tau+1}}\left(I(\tau)-I(\tau+1)\right)-\frac{C_{\tau+1}}{S_{\tau+1}}\Delta_{\tau},\end{split}$ (48)
which also means that for $I(\tau)$ strictly decreasing, $\Delta_{\tau}$ is initially increasing driven by the first term $S_{\tau}\left(I(\tau)-I(\tau+1)\right)/S_{\tau+1}$. Then it starts decreasing for $S_{\tau}\left(I(\tau)-I(\tau+1)\right)<C_{\tau+1}\Delta_{\tau}$. In the deterministic limit, we approximate $C_{\tau}\sim e^{r\tau}$ where $r$ is the growth rate, and obtain that this condition is
$\Delta_{\tau}>\left(I(\tau)-I(\tau+1)\right)\frac{1-e^{-r(\tau+1)}}{e^{r}-1}.$ (49)
For a decreasing yearly intensity, $\Delta_{\tau}$ first increases until this condition is reached, then evolves depending on the sign of $\left(I(\tau)-I(\tau+1)\right)$. Then at large times, we get
$\Delta_{\tau+1}\sim\left(I(\tau)-I(\tau+1)+\Delta_{\tau}\right)e^{-r}.$ (50)
From this, we deduce that at large times
$\Delta_{\tau}\sim-\frac{1}{e^{r}-1}\frac{dI}{d\tau},$ (51)
meaning that $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\sim I(\tau)-(dI/d\tau)/(e^{r}-1)$. In particular if we assume a power law dependency for the carbon intensity of the form $I(\tau)=I_{0}-\eta\tau^{n}$, the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ goes as
$\Delta_{\tau}\sim\frac{\eta n\tau^{n-1}}{e^{r}-1},$ (52)
therefore increasing to $\infty$ for $n>1$ and decreasing to 0 for $n<1$. In practice it means that for fast decarbonization, early intense years matter more in the path dependent intensity.
Illustration with linearly decreasing intensities
If we use a linear relationship $I(\tau)=I_{0}-\eta(\tau-\tau_{0})$ (which seems to be observed in most cases [38]), we get that
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)=I_{0}-\eta\frac{\sum_{\tau^{\prime}=0}^{\tau}\tau^{\prime}C_{\tau^{\prime}}}{\sum_{\tau^{\prime}=0}^{\tau}C_{\tau^{\prime}}},$ (53)
In addition, if we forget about stochasticity and assume exponential growth $C_{\tau}=LC_{\text{ini}}e^{r\tau}$, we get
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)=I_{0}-\frac{\eta e^{r}}{e^{r}-1}\left(\frac{e^{r\tau}(\tau+1)(e^{r}-1)}{e^{r(\tau+1)}-1}-1\right),$ (54)
and finally
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)=\eta\left(\tau+\frac{e^{r}((\tau+1)e^{r\tau}-\tau e^{r(\tau+1)}-1))}{(e^{r}-1)(e^{r(\tau+1)}-1)}\right),$ (55)
which is positive and increasing towards $1/(e^{r}-1)$ provided that $r>0$ as can be seen on Fig.9. This means that path dependent intensities are always larger than yearly intensities whenever GDP is growing over time.
Figure 9: A Difference between the path dependent intensity $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and the yearly intensity $I(\tau)$ with a deterministic evolution of $C_{\tau}$. The path-dependent intensity is systematically larger than the yearly intensity for $r>0$ and increases with time. It eventually reaches a strictly positive limit. B Effect of stochasticity on the difference between the path-dependent intensity $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and the yearly intensity $I(\tau)$. Increasing fluctuations modify the difference between the two quantities. C $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)$ follows a power law $n\eta\tau^{n-1}/(e^{r}-1)$ after a first transient phase, depending on the decarbonization rate.
With stochasticity
If we now add stochasticity and use $\pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right):=\prod_{1\leq\tau^{\prime}\leq\tau}V_{\tau^{\prime}}$, we have $C_{\tau}=LC_{\text{ini}}\pi(\tau)$. In addition, with $X_{\tau^{\prime}}=\ln(V_{\tau^{\prime}})$, we have that $C_{\tau}=LC_{\text{ini}}\prod_{1\leq\tau^{\prime}\leq\tau}e^{X_{\tau^{\prime}}}$. In the limit where $X_{\tau^{\prime}}$ is drawn according to a normal distribution with mean $r$ and standard deviation $\sigma_{X}$, we write $X_{\tau^{\prime}}=r+\sigma_{X}\xi_{\tau^{\prime}}$ ($\xi_{\tau^{\prime}}$ is normally distributed with mean 0 and standard deviation $1$). We can thus write
$\begin{split}\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)&=\frac{\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau)\right)e^{r\tau^{\prime}}\prod_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}e^{\sigma_{X}\xi_{\tau^{\prime\prime}}}}{\sum_{\tau^{\prime}=0}^{\tau}e^{r\tau^{\prime}}\prod_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}e^{\sigma_{X}\xi_{\tau^{\prime\prime}}}}\\ &=\frac{\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau)\right)\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)}{\sum_{\tau^{\prime}=0}^{\tau}\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)}.\end{split}$ (56)
Since $\xi(\tau)$ is normally distributed, we get that $\left<\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)\right>=\exp\left((r+\sigma_{X}^{2}/2)\tau^{\prime}\right)$, and
$\mathbb{V}(\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right))=\exp\left((2r+\sigma_{X}^{2})\tau^{\prime}\right)\left(\exp\left(\sigma_{X}^{2}\tau^{\prime}\right)-1\right).$ (57)
Therefore this variable has wide fluctuations. For small $\sigma_{X}$, we recover the previous deterministic result. We show the dependency of this on the fluctuations $\sigma_{X}$ on Fig.9B. We observe that the difference between path dependent intensity and yearly intensity increases with fluctuations $\sigma_{X}$. This results from the fact that if a phase of large economic growth with large carbon intensities is followed by a phase of recessions with lower carbon intensities (years with a stochastically low growth rate), the weight of carbon-emissive years will be larger. Therefore, the weights of years with large carbon intensities will be larger and the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ will increase, leading to a larger decoupling. In addition, large fluctuations may lead to years with a stochastically large growth rate and low carbon intensities, typically reducing the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$. However, as shown in the deterministic case, the difference typically goes as $1/r$, so that decreasing the growth rate stochastically has a larger impact on decoupling than increasing the growth rate stochastically does. Therefore, recessions lead to an increased decoupling between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ and the average decoupling is larger when fluctuations are larger.
To better understand this effect we can study what happens for a first phase with a high growth rate $\gamma_{h}$, followed by a phase of recession with a lower growth rate $\gamma_{l}$. In this case, for a phase of fast growth lasting for $\tau_{h}$ and a phase of recession lasting for $\tau-\tau_{h}$, we get
$\begin{split}\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)&=\frac{(e^{\gamma_{h}}-1)(e^{\gamma_{l}}-1)}{(e^{\gamma_{l}}-1)(e^{\gamma_{l}(\tau_{h}+1)}-1)+(e^{\gamma_{h}}-1)(e^{\gamma_{h}(\tau+1)}-e^{\gamma_{h}(\tau_{h}+1)})}\times\\ &\left(\frac{\left((e^{\gamma_{h}}-1)((\tau-\tau_{l})e^{\gamma_{h}(\tau_{h}+1)}-\tau)-e^{\gamma_{h}}(1+e^{\gamma_{h}\tau_{h}})\right)}{(e^{\gamma_{h}}-1)^{2}}+\frac{\left((e^{\gamma_{l}}-1)(\tau_{l}-\tau)e^{\gamma_{l}(\tau_{h}+1)}-e^{\gamma_{l}(\tau+1)}+e^{\gamma_{l}(\tau_{h}+1)}\right)}{(e^{\gamma_{l}}-1)^{2}}\right),\end{split}$ (58)
So that for large $\tau$, it scales as the duration of the recession divided by $e^{\lambda_{l}}-1$
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)\sim\frac{\tau-\tau_{h}}{e^{\gamma_{l}}-1}$ (59)
Therefore, only the growth rate and duration of the recession phase matter to explain the decoupling between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ in this limit.
B.4 Study of the solvency probability
Strategy A
We can eventually relate this to the probability that a biased random walk is smaller than a threshold at payback time $t_{p}$ [52, 25] (the jumps do not have a zero mean, and the random walk is thus biased). We emphasize that bankruptcy does not correspond to the existence of a time where the debt is larger than the gross capital (or first passage time), but to the fact that at the time of payback $t_{p}$, the debt is larger than the gross capital.
By defining $C_{\tau^{\prime}}=\ln(\gamma_{\tau^{\prime}}/\rho)$ and $S_{\tau}=\sum_{\tau^{\prime}\leq\tau}C_{\tau^{\prime}}$. We assume that the asymetrical distribution of $X$ is $\phi$, with finite mean $m=W-\ln(\rho)$ and variance $\sigma_{X}^{2}$. Using the central limit theorem for large enough $\tau$, we finally obtain that
$\begin{split}P_{S}^{A}(t_{p})&=\mathbb{P}\left(\sum_{\tau^{\prime}\leq\tau}C_{\tau^{\prime}}\geq\ln\left(1-\frac{1}{L}\right)\right)\\ &=\frac{1}{2}\left(1+\text{erf}\left(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}}\right)\right)\end{split}$ (60)
We can study the argument of the error function and how it depends on time. In particular, if $m>0$ ( $W>\ln(\rho)$), it is a non-monotonic function, decreasing until $t_{p}=-\ln(1-1/L)/(W-\ln(\rho))$ (worst payback time) and then increasing towards $1$. Otherwise, it is a strictly decreasing function of time. Therefore, the result will strongly depend on the sign of $m$.
Figure 10: Solvency probability with Gaussian jumps, but the same average and standard deviation as in Kelly’s case in the limit of large leverage $L=10$ with strategy A.
We plot the solvency probability as a function of the payback time $t_{p}$ in Fig.10 in the limit of large leverage. The short-term growth rate $w(t_{p})$ for strategy A is shown in Fig.11. Using the expression of the short-term growth rate with strategy A, we have that this is defined for all times if $E\geq\rho$. Otherwise this is well defined for all times lower than $\ln(1-1/L)/(\ln(\rho)-\ln(E))$. In addition, if we define
$g(t_{p})=E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}},$ (61)
which is the argument of the $\ln$ in the definition of the short-term growth rate, we have that
$w^{\prime}(t_{p})=\frac{1}{t_{p}^{2}}\left(t_{p}\frac{g^{\prime}(t_{p})}{g(t_{p})}-\ln(g(t_{p}))\right).$ (62)
The latter is equal to 0 if and only if $t_{p}g^{\prime}(t_{p})/g(t_{p})=\ln(g(t_{p}))$. Concerning the solvency probability, we observe that this is indeed a decreasing function of the leverage (higher leverage is associated with higher risk). We thus identify four regimes based on the concavity inequality $\ln(E)\geq W$ :
  • If $m>0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 1$, faster as $\sigma_{X}$ decreases. The solvency probability is a non monotonic function of $t_{p}$, first decreasing until $t_{p}=-\ln(1-1/L)/(W-\ln(\rho))$, then increasing towards $1$. For large leverage the first regime vanishes (when $\ln(1-1/L)/(W-\ln(\rho))<1$), but it means that for small enough leverage it may be safer to pay back as soon as possible, or to wait for long enough. The short-term growth rate is well defined at all times.
  • If $m=0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)=\frac{1}{2}$. In this case the solvency probability is decreasing with $t_{p}$. The short term growth rate is well defined at all times.
  • If $m<0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. The solvency probability is decreasing with $t_{p}$. The short term growth rate is well defined at all times.
  • If $m<0$ and $\ln(E)<\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. The solvency probability is decreasing with $t_{p}$. The short term growth rate is only defined for $t_{p}<-\ln(1-1/L)/(\ln(\rho)-\ln(E))$.
Those results are summed up on Fig.11.
Figure 11: solvency probability and short-term growth rate for strategy A, different values of leverage and interest rates. For the solvency probability, there is a transition at $\ln(\rho)=W$. When $W<\ln(\rho)$, the solvency probability is strictly decreasing for all times; otherwise, it is a non-monotonic function of time, with a minimum at $t_{p}=-\ln(1-1/L)/(W-\ln(\rho))$. For the short-term growth rate, there is a transition at $\rho=E$. When the interest rate increases above $E$ and remains below $E^{L/(L-1)}$, the short term growth rate starts decreasing with time after $t_{p}>\ln((1-1/L)\ln(\rho)/\ln(E))/\ln(E/\rho)$. The figure on the left corresponds to $\ln(\rho)<W$. The figure on the right corresponds to $W<\ln(\rho)$
Strategy B
Using strategy B, the solvency probability is slightly different, and we can’t used the central limit theorem anymore
$\begin{split}P_{S}^{B}(t_{p})&=\mathbb{P}\left(\prod_{1\leq\tau^{\prime}\leq t_{p}}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)-\sum_{0\leq t\leq t_{p}-1}\frac{L-1}{\rho}\prod_{1\leq\tau^{\prime}\leq t}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\geq 1-\frac{1}{L}\right).\end{split}$ (63)
This suggests to introduce $\Pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)/\rho$, so that
$\begin{split}P_{S}^{B}(t_{p})&=\mathbb{P}\left(\Pi_{t_{p}}-\frac{L-1}{\rho}\sum_{0\leq t\leq t_{p}-1}\Pi_{t}\geq 1-\frac{1}{L}\right),\end{split}$ (64)
and
$\begin{split}C_{t_{p}}&=LC_{\text{ini}}\rho^{t_{p}}\Pi_{t_{p}}\\ D_{t_{p}}&=LC_{\text{ini}}\rho^{t_{p}}\left(1-\frac{1}{L}+\frac{L-1}{\rho}\sum_{0\leq t\leq t_{p}-1}\Pi_{t}\right).\end{split}$ (65)
Now we can write
$\Pi(\tau)=\exp\left(\sum_{1\leq\tau^{\prime}\leq\tau}\ln(V_{\tau^{\prime}})\right)=\prod_{1\leq\tau^{\prime}\leq\tau}V_{\tau^{\prime}},$ (66)
and using the Central limit theorem, $\sum_{\tau^{\prime}\leq\tau}\ln(V_{\tau^{\prime}})$ follows a normal law $\mathcal{N}\left(\tau\mu,\sqrt{\tau}\sigma_{X}\right)$. Therefore $\Pi(\tau)$ is log-normally distributed, with mean $\exp\left(\tau(\mu+\sigma_{X}^{2}/2)\right)$ and variance $(\exp(\tau\sigma_{X}^{2})-1)\exp(\tau(2\mu+\sigma_{X}^{2}))$. Its probability distribution function is:
$\pi(\tau)(x)=\frac{1}{\sqrt{2\pi\tau}x\sigma_{X}}\exp\left(-\frac{(\ln(x)-\mu\tau)^{2}}{2\tau\sigma_{X}^{2}}\right).$ (67)
If we define the quantity (in terms of the capital $C_{t_{p}}$ and debt $D_{t_{p}}$ at time $t_{p}$)
$\begin{split}X_{t_{p}}&=\frac{C_{t_{p}}-D_{t_{p}}}{LC_{\text{ini}}\rho^{t_{p}}}+1-\frac{1}{L}\\ &=\Pi_{t_{p}}-\frac{L-1}{\rho}\sum_{0\leq t\leq t_{p}-1}\Pi_{t}\\ &=V_{1}...V_{t_{p}}-\frac{L-1}{\rho}\left(1+V_{1}+V_{1}V_{2}+...+V_{1}...V_{t_{p}-1}\right)\\ &=-\frac{L-1}{\rho}+V_{1}\left(V_{2}...V_{t_{p}}-\frac{L-1}{\rho}\left(1+V_{2}+V_{2}V_{3}+...+V_{2}...V_{t_{p}-1}\right)\right)\\ &=-\frac{L-1}{\rho}+V_{1}\left(-\frac{L-1}{\rho}+V_{2}\left(V_{3}...V_{t_{p}}-\frac{L-1}{\rho}\left(1+V_{3}+V_{3}V_{4}+...+V_{3}...V_{t_{p}-1}\right)\right)\right)\\ &=-\frac{L-1}{\rho}+V_{1}\left(-\frac{L-1}{\rho}+V_{2}\left(-\frac{L-1}{\rho}+V_{3}\left(-\frac{L-1}{\rho}+...(-\frac{L-1}{\rho}+V_{t_{p}})\right)\right)\right),\end{split}$ (68)
The sign of $X_{t_{p}}-(1-1/L)$ encapsulates solvency at time $t_{p}$. We can study the monotonicity of $X_{t_{p}}$, in particular we have that
$\begin{split}X_{t_{p}+1}-X_{t_{p}}&=(V_{tp+1}-1)V_{0}...V_{t_{p}}-\frac{L-1}{\rho}V_{0}...V_{t_{p}}\\ &=V_{0}V_{1}...V_{t_{p}}\left(V_{t_{p}+1}-\left(1+\frac{L-1}{\rho}\right)\right)\\ &=V_{0}V_{1}...V_{t_{p}}\left(\frac{\gamma_{t_{p}+1}}{\rho}-1\right).\end{split}$ (69)
Therefore, using that the $V_{\tau}$ and $\gamma_{\tau}$ are independent, we get that in average the sign of $X_{t_{p}+1}-X_{t_{p}}$ is set by the sign of $\left<\gamma\right>-\rho$. We can thus draw two conclusions
  • If $\left<\gamma\right><\rho$ (which seems to be the case in most countries), we have that if the debt is larger than the capital at time $t_{p}$ then it is also the case at time $t_{p}+1$ in average. For a fully deterministic model it would mean that the survival probability is a decreasing function of leverage. However due to stochasticity, this result is no longer true when $\gamma$ has a broad distribution (the distribution is broader for higher $\lambda$).
    $\begin{split}\left(\left<X_{t_{p}+1}\right><1-\frac{1}{L}\right)\subset\left(\left<X_{t_{p}}\right><1-\frac{1}{L}\right)\end{split}$ (70)
  • Else if $\left<\gamma\right>>\rho$, we have that if the debt is lower than the capital at time $t_{p}$ then it is also the case at time $t_{p}+1$ in average
    $\left(\left<X_{t_{p}+1}\right>>1-\frac{1}{L}\right)\subset\left(\left<X_{t_{p}}\right>>1-\frac{1}{L}\right)$ (71)
In addition, the solvency probability is related to the cumulative distribution function of a sum of correlated log-normally distributed variables. If we assume that the jumps are normally distributed, $V_{\tau}$ are log-normally distributed and independent. We introduce the probability distribution function of a log-normal distribution
$\phi(x)=\frac{1}{\sqrt{2\pi}x\sigma_{X}}\exp\left(-\frac{(\ln(x)-\mu)^{2}}{2\sigma_{X}^{2}}\right).$ (72)
We plot it on Fig.12, where it reaches a plateau for small enough interest rates. We see that solvency is smaller with strategy B, a risk that has to be taken to sustain a larger growth rate. For $L\sim 1$ we understand why this solvency probability (strategy B) is smaller than for strategy A, as the lower bound for solvency is larger.
Figure 12: Comparing strategies with the same leverage ($L=1.05$), as expected, strategy A has a higher solvency probability. Yet it also leads to a lower capital in the end.
B.5 Optimal leverage and carbon emissions
For strategy B, we get that the average growth rate is increasing with the leverage $\sim\left<\log\left(\gamma_{t}+L-1\right)\right$>, while the survival probability has a non monotonous complex behaviour as a function of leverage. We can build an optimization function between growth rate and the survival probability
$J(\beta,L,t_{p})=\beta\left<\log\left(\gamma_{t}+L-1\right)\right>+(1-\beta)P_{S}^{B}(t_{p})$ (73)
Typically for the range of values corresponding to real data, we obtain the behaviour of Fig. 13. While the survival probability first decreases with leverage then plateaus or increases slowly, the growth rate keeps increasing with leverage. Therefore, the optimization function first decreases, reaches a minimum, and then increases again (the increase in growth rate compensates the plateau in survival probability). From this, we deduce that high leverages could be considered as optimal. However this optimization function does not account for carbon emissions resulting from a stronger growth rate. This suggests that the ”optimal” strategy with credit in a carbon intensive economy leads to stronger cumulative emissions than without credit.
Figure 13: Optimization function for the leverage with $\beta=0.8,\rho=2,\left<\log(\gamma)\right>=0.7$ and for different values of $t_{p}$.
B.6 Optimal payback time based on solvency and short term growth rate
Let us start by reminding the key parameters in this version of the model for solvency:
  • $W$ is the long-term average growth rate, provided there was no bankruptcy; otherwise, this is 0.
  • $P_{S}(t_{p})$ is the probability of solvency after payback $t_{p}$.
  • $\sigma_{X}$ is the standard deviation of the long-term growth rate. It encapsulates the fluctuations of the system. We highlight that this is also the standard deviation of $m=W-\ln(\rho)$.
  • $WP_{S}(t_{p})$ is the effective long-term growth rate, conditioned on solvency after pay-back (otherwise the growth rate is 0).
  • The average growth rate of the capital relative to the initial investment $LC_{\text{ini}}$ just after the debt has been paid back, given there was no bankruptcy, $w(t_{p})$. It represents a short term growth rate.
The gambler wants to maximize the average growth rate $W$, the solvency probability $P_{S}(t_{p})$, minimize fluctuations $\sigma_{X}$ and maximize the average relative earnings after payback, keeping in mind that those parameters are interdependent.
We can first form the same optimization function as in [23]:
$J_{1}(\alpha)=\alpha W-(1-\alpha)\sigma_{X},$ (74)
where the value $\alpha$ sets the strategy. For $\alpha=1$ fluctuations (proxy for the risk) are ignored and only the growth rate is maximized, which corresponds to Kelly’s optimal strategy, while for $\alpha=0$ only fluctuations are minimized. Safe strategies correspond the strategies where fluctuations are reduced, even at the cost of reducing the growth rate.
If we optimize this with respect to the betting strategy $b_{x}$, under the constraint that $\sum_{i}b_{i}=1$ (Lagrange multiplier $\mu$), we get:
$\frac{\partial J_{1}}{\partial b_{i}}=$ $\alpha\frac{p_{i}}{b_{i}}-(1-\alpha)\frac{p_{i}}{\sigma_{X}b_{i}}\left(\ln(\frac{b_{i}}{r_{i}})-W\right)-\mu$ (75)
We deduce that:
$\mu b_{i}=\alpha p_{i}-\frac{1-\alpha}{\sigma_{X}}p_{i}\left(\ln\left(\frac{b_{i}}{r_{i}}\right)-W\right)$ (76)
Therefore, imposing that $\sum_{i}b_{i}=1$, we get a set of equations:
$\mu$ $=\alpha$ (77)
$b_{i}$ $=p_{i}\left(1-\frac{1-\alpha}{\alpha\sigma_{X}}\left(\ln\left(\frac{b_{i}}{r_{i}}\right)-W\right)\right)$ (78)
Thus we recover a correction from Kelly’s optimum $b_{i}=p_{i}$, depending on $\alpha,\sigma_{X}$ and $\boldsymbol{r}$. In this case, the bet should be modified particularly for horses corresponding to a value of $\ln(b_{i}/r_{i})$ significantly different from their average $W$. In particular, for a distribution of $b_{i}/r_{i}$ with fat tails.
The payback time is a balance between the need to accumulate as much as possible before payback and the risk of waiting too long before paying back, particularly with high interest rates. Indeed, if the interest rate is significantly smaller than the average earnings, the best strategy is to wait as long as possible to repay. Yet, if the interest rate is close to the average earnings, waiting too long may become risky due to fluctuations or simply because debts accumulate.
Strategy A
We observe that the Gaussian approximation seems to describe accurately the discrete case. Therefore we propose to use this expression of $P_{S}(t_{p})$ to optimize payback time with respect to solvency probability.
On another hand we can optimize the short term growth rate $w(t_{p})$ with respect to payback time $t_{p}$.
in a second optimization function with a new objective function $J_{2}(\beta)$, that interpolates between maximizing the solvency probability and maximizing short term growth rate relative to $\ln(E)$. We find
$J_{2}(\beta)=\beta\frac{1}{t_{p}\ln(E)}\ln\left(E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}\right)+\frac{1}{2}(1-\beta)\left(1+\text{erf}(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}})\right).$ (79)
The landscape of this optimization function is shown in Fig.14. Then, using $m=W-\ln(\rho)$, we are able to express:
$\begin{split}\frac{\partial J_{2}}{\partial t_{p}}=&\frac{\beta}{t_{p}^{2}\ln(E)}\left[t_{p}\frac{\ln(E)E^{t_{p}}-(1-\frac{1}{L})\ln(\rho)\rho^{t_{p}}}{E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}}-\ln\left(E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}\right)\right]\\ &+(1-\beta)\frac{W-\ln(\rho)+\frac{\ln(1-\frac{1}{L})}{t_{p}}}{2\sigma_{X}\sqrt{2\pi t_{p}}}\exp\left(-\frac{\left(t_{p}(W-\ln(\rho))-\ln\left(1-\frac{1}{L}\right)\right)^{2}}{2t_{p}\sigma_{X}^{2}}\right)\end{split}$ (80)
Figure 14: Landscape of the optimization function for two different sets of horses with $\beta=0.5$ in the limit of large leverage $L=15$.
Let us start with the limit of large $L$, we have
$\begin{split}\frac{\partial J_{2}}{\partial t_{p}}=&\frac{\beta}{t_{p}^{2}\ln(E)}\left[t_{p}\frac{\ln(E)E^{t_{p}}-\ln(\rho)\rho^{t_{p}}}{E^{t_{p}}-\rho^{t_{p}}}-\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)\right]+(1-\beta)\frac{W-\ln(\rho)}{2\sigma_{X}\sqrt{2\pi t_{p}}}\exp\left(t_{p}\frac{\left(W-\ln(\rho)\right)^{2}}{2\sigma_{X}^{2}}\right)\end{split}$ (81)
We remind that the concavity of $\ln$ ensures that $W\leq\ln(E)$. We can compare three regions of the space $\left(\ln(\rho),\ln(E)\right)$ : $\ln(E)>W>\ln(\rho)$, $\ln(E)>\ln(\rho)>W$ and $\ln(\rho)>\ln(E)>W$ in the limit of large $L$
  • If $\ln(E)>W>\ln(\rho)$:
    $\frac{t_{p}^{3/2}(E^{t_{p}}-\rho^{t_{p}})\exp\left(\left(\left(\frac{(W-\ln(\rho))^{2}}{2\sigma_{X}^{2}}\right)-\ln(\frac{\rho}{E})\right)t_{p}\right)}{(\ln(E^{t_{p}})E^{t_{p}}-\ln(\rho^{t_{p}})\rho^{t_{p}})-(E^{t_{p}}-\rho^{t_{p}})\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)}$ $=-\frac{\beta}{(1-\beta)\ln(E)}\frac{2\sqrt{2\pi}\sigma_{X}}{W-\ln(\rho)}<0,$ (82)
    this has no finite solution. Indeed we can show that $x\mapsto(x\ln(x)-\rho^{t_{p}}\ln(\rho^{t_{p}}))-(x-\rho^{t_{p}})\ln(x-\rho^{t_{p}})$ is an increasing function and thus $(\ln(E^{t_{p}})E^{t_{p}}-\ln(\rho^{t_{p}})\rho^{t_{p}})-(E^{t_{p}}-\rho^{t_{p}})\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)\geq 0$. Considering that $J_{2}(\alpha,p_{x},b_{x},r_{x},t_{p})$ is increasing with $t_{p}$, it will be maximum for $t_{p}\to\infty$.
  • If $\ln(E)>\ln(\rho)>W$, then:
    $\frac{t_{p}^{3/2}(E^{t_{p}}-\rho^{t_{p}})\exp\left(\left(\left(\frac{(W-\ln(\rho))^{2}}{2\sigma_{X}^{2}}\right)-\ln(\frac{\rho}{E})\right)t_{p}\right)}{(\ln(E^{t_{p}})E^{t_{p}}-\ln(\rho^{t_{p}})\rho^{t_{p}})-(E^{t_{p}}-\rho^{t_{p}})\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)}$ $=-\frac{\beta}{(1-\beta)\ln(E)}\frac{2\sqrt{2\pi}\sigma_{X}}{W-\ln(\rho)}>0,$ (83)
    this has finite solutions a priori.
  • If $\ln(\rho)>\ln(E)>W$, $J_{2}$, $w(t_{p})$ is not defined, we thus consider that the optimality is only set by maximizing the solvency probability. The optimum is thus $t_{p}\to 0$.
Now, when $L$ is small, new inequalities become decisive
$\begin{split}&\ln(\rho)<\ln(E)+\frac{1}{t_{p}}\ln(1+\frac{1}{L})\\ &\ln(\rho)<W+\frac{1}{t_{p}}\ln(1-\frac{1}{L}),\end{split}$ (84)
The first is the condition to have a well-defined short-term growth rate, and the second sets the transition to a solvency probability increasing with $t_{p}$ and a solvency probability decreasing with $t_{p}$. The first inequality means that the short-term growth rate can be defined even for $\ln(\rho)>\ln(E)$ (at small $t_{p}$). The second inequality means that the solvency probability can be a decreasing function of $t_{p}$ even for $W>\ln(\rho)$.
Figure 15: Optimal payback time depending on the conditions and the importance given to solvency probability. We observe that when we increase $\beta$, we grant less importance to solvency and waiting for longer becomes favorable. In addition, when increasing leverage $L$, we increase the short term growth rate, and we decrease solvency probability. Yet solvency probability becomes considerably smaller for large times, while the difference is smaller at early times. Therefore, increasing the leverage at fixed $\beta$ tends to increase optimal payback time. The region $E<\rho$ is never favorable.
We see on Fig.15 that the optimal payback time depends on $\rho$, $E$, $L$ and $\beta$. $\beta$ quantifies the relative importance given to short growth rate versus solvency probability. As $\beta$ increases, the gambler becomes less careful and wants to optimize the short-term growth rate rather than the solvency probability; in this case, the optimal payback time increases. As $L$ increases, the short-term growth rate also increases, and the solvency probability typically decreases, yet more significantly at long times. This tends to increase the optimal payback time. There are two main transitions in all diagrams, located around $E=\rho$ and $W=\ln(\rho)$ (shifted for small leverages), corresponding respectively to the transition where short-term growth rate starts being undefined at long times and to the transition where solvency probability starts decreasing with time.
B.7 Growth rate-solvency probability trade-off
The long-term growth rate is $W$ if the gambler survives to payback, and 0 otherwise. We can thus introduce an effective growth rate $WP_{S}(t_{p})$. We obtain the different dependencies of the parameters with respect to one another on Fig.16. We recover the results of [23] on Fig.16, where there is a trade-off branch where $\sigma_{X}$ is increasing with $WP_{S}(t_{p})$ (the effective growth rate). Kelly’s strategy corresponds to the maximal effective growth rate. We see on Fig.16 that $P_{S}(t_{p})$ is decreasing with $\sigma_{X}$ (increasing fluctuations means decreasing the probability of solvency), and that $w(t_{p})$ is increasing with $WP_{S}(t_{p})$ along the safe trajectory. Finally, we observe another trade-off in Fig.16, where $P_{S}(t_{p})$ decreases with $w(t_{p})$, meaning that ensuring solvency is done at the cost of reducing the short-term growth rate.
Figure 16: Different trade-offs between key parameters. In particular, increasing the effective growth rate goes along with increasing fluctuations. We also observe that increasing solvency probability goes along with decreasing short-term growth rate. Kelly’s strategy is shown as a square for each set of parameters, and safe strategies (varying $\alpha$) are shown as straight lines.
We highlight that for the safe strategy (full lines), fluctuations are smaller, the solvency probability is usually larger(except for large interest rates), the growth rate is smaller, and the post-payback earnings are smaller. The safe strategy corresponds to the optimization done in [23], with the parameter $\alpha$ varying.
Appendix C Data
All data used in this work are publicly available for the different countries under study. In particular we get data for the US in [43, 45, 42, 44]. For France, we use [77, 78, 79, 80]. For Denmark, we use [81, 82, 83, 84]. For China, we use [85, 86, 87, 88].